Shift summary for 8/16/99 DAY shift

Julie Whitmore (jaws@fnpxsr)
Mon, 16 Aug 1999 15:29:26 CDT

Shift Summary for DAY shift, 8/16/99

On Shift: Erik R., Julie

Hours of beam available at all intensities: 7.5
Hours of beam available at nominal intensity: 7.25
Hours of E832 data (special runs not included): 6.5
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: ktev1 UPSII error, kfrends, accelerator work
on F3 wet engine.

========================================================================
Runs ended this shift:
Run 14236 - 23.8M events - BEAM_832
Run 14237 - 15.9M events - BEAM_832

Tapes brought to FCC this shift:
(including tapes for online split)

QKC270,QKC271,QKC272
QKK270,QKK271,QKK272
QKS270,QKS271,QKS272
QKB096

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:
Install calibration constants for CsI. Boost quiescent laser light.
Smooth E832 data taking under Sydney's benign rule.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
07:01 Julie removes her name as run manager from
white board.

07:05 Sydney arrives to take the helm.

14236 07:35 Run ends. Tapes full. DRT_FRESH_START to
recover L3 process and online displays.

07:40 DRT_FRESH_START fails with ktev1 timeout:
fails to initialize UPSII environment. Sydney
calls Etta. The problem is a large core file
(presumably from the ktev1 death on Sun owl shift)
which takes up space on ktev1's root partition.
She cleans this up. Fresh start now works.

14237 1-3 08:10 Start run. HCC bits missing for TW quadrant.
Reloading Et maps clears it.

85-91 10:03 ktevana errors and HCC problem.
Reload maps for HCC and Et. kfrend.

93-94 10:16 Lose HCC bits for TW quadrant. Reload Et.

10:40 Doug: having trouble with data logging on
vacuum PC. Power cycle PC and restarting FIX
a couple of times, dta logging is now fine.

140 11:20 MCR notifies us that they are dropping back to
4E12 (1 turn) to study injection losses.

150 11:30 Laser RMS wide from channel 1137. OK after
1 spill. Quiescent laser shows that channel is
still having HV problems.

153 11:33 Intensity back up to ~6E12.

178-187 12:10 ktevana errors. Needed to cycle laser crate
repeater. kfrend now clears.

12:30 ECC remakes L3_driver_mon_832 to fix bug that
causes ktev3 process to crash.

218 13:00 ktevana errors. kfrend.

258-261 13:57 ktevana errors. Multiple kfrends+pipeline reset.

14:00 Beam disappears. "Return F3 wet engine to
service. Estimate for beam 20min."

14:10 Run ended. PeterSh increases quiescent
laser light. EdB installs new CsI calibration
constants and updates PMT gain drift reference
plot. Brute force L3 to pick up new exectable.

14238 14:23 Start run.

26 14:55 Beam disappears. "MI power supply problem.
No estimate for beam."

========================================================================
Follow up Notes on Previous Shift
Dart fresh start done.

========================================================================
Items to Follow up on Next Shift
Label QKB**** tapes.